OpenStudy (anonymous):

Solve for x from this log equation (Shown in next post) I think I did it wrong, as I couldn't isolate x on one side.

OpenStudy (anonymous):

\[1.32(\ln ^{0.4}(4.65x ^{2}))x ^{0.8}=s\] The furthest I got was this: \[x ^{2}\ln(x)+0.76935x ^{2}= \frac{ s }{ 2.64 }^{(5/2)}\]

OpenStudy (anonymous):

That gives me ln(x^(x^2))..., which I can't do anything with

OpenStudy (anonymous):

i though i had something but i don't you have x in the log and outside the log right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wolfram talks about...complex logs >_>

OpenStudy (anonymous):

yeah i am pretty sure you cannot separate them

OpenStudy (anonymous):

Maybe I did something wrong, and broke a rule somewhere

OpenStudy (anonymous):

is that the original question?

OpenStudy (anonymous):

what is? No, I *think* I did, because I can't get x on one side

OpenStudy (anonymous):

i meant is \[1.32\ln ^{0.4}(4.65x ^{2})x ^{0.8}=s\] the original ?

OpenStudy (anonymous):

Yes that's the original equation.

OpenStudy (anonymous):

pretty sure it can;t be done you cannot even solve for x in \[x\ln(x)=s\]

OpenStudy (anonymous):

Did I at least get my furthest equation right?

OpenStudy (anonymous):

*as simplified as I did

OpenStudy (anonymous):

i honestly can't tell what you did, but i don't think so

OpenStudy (anonymous):

I can show the steps, but if it's unsolavable it might as well not matter

OpenStudy (anonymous):

Thanks for the help

OpenStudy (anonymous):

yeah don't fret about this not worth it sorry

OpenStudy (anonymous):

yw